# Huge Ensembles Part II: Properties of a Huge Ensemble of Hindcasts Generated with Spherical Fourier Neural Operators

### v2

@@ 4.2 Confidence Intervals of Extreme Forecasts
In summer 2023, over all 10-day forecasts at all grid cells, the confidence intervals are an order of magnitude narrower than those from small ensembles (Figure 10). This result is robust across different lead times: Figure 15 calculates the same confidence interval width at lead times of 4 days and 7 days. With less sampling uncertainty for extreme forecasts, there could be more informed disaster readiness and more targeted plans. HENS shrinks the confidence intervals because it thoroughly represents the forecast distribution and reduces sampling uncertainty. However, even with narrower confidence intervals, we note that it is still of course possible to have uncertain forecasts. For instance, if half the ensemble members predict extreme weather, the chance of extreme would still be 50%. With HENS, the 95 percent confidence interval around the 50% estimate would be narrower.
@@ 4.3 Missed Events in HENS and IFS
The outlier statistic measures the ability of an ensemble to capture a true event (Buizza and Palmer, 1998). The outlier statistic measures the proportion of cases in which an observed event lies outside the bounds of the ensemble forecast. Using HENS, we calculate this statistic as a function of ensemble size. For each ensemble size, we take 100 bootstrap random samples from HENS. At each grid cell, if the ERA5 value lies within ninety-five percent of the bootstrapped ensembles, then the ensemble size is deemed satisfactory. Otherwise, the ERA5 value is classified as an outlier. The outlier statistic is the proportion of grid cells that are outliers for a given ensemble size. By taking 100 bootstrap random samples, we consider sampling uncertainty in the outlier statistic.
Figure 11 shows larger ensembles reduce the probability of an outlier. At a lead time of 10 days, 7,424 member ensemble is satisfactory for representing almost 99% of the ERA5 values. The reduction in the outlier statistic is robust across lead times of 4 days, 7 days, and 10 days. Larger ensembles can greatly increase the coverage of all the possible solutions of how the weather at each grid cell can evolve. Regardless of ensemble size, we note that 10 day forecasts have a lower outlier statistic than 7-day and 4-day forecasts because they have the best spread-error ratio (see Part 1).
By taking 100 bootstrap random samples, we modify the calculation of the original statistic presented in Buizza and Palmer (1998). Buizza and Palmer (1998) only consider one possible ensemble for each ensemble size, so they do not characterize sampling uncertainty in the outlier statistic. With bootstrap samples, it is easier for a grid cell to be an outlier. Therefore, the outlier statistic increases, since a given ensemble size cannot simply capture the event with just one member. It must capture the event in 95 percent of bootstrap random samples. For a more direct comparison with the Buizza and Palmer (1998) method, we also calculate the proportion of cases in which the verification dataset is greater than the ensemble max. We do not consider ensemble sampling uncertainty, and this is analogous to the warm side end of their outlier statistic. We choose to focus only on the warm end because summer 2023 was the hottest summer on record, and we particularly assess the ability of HENS to represent extreme heatwaves.
@@
Across all of summer 2023, Figure 12a shows that the bulk of the distribution lies above the 1-to-1 line, indicating that the HENS maximum successfully captures the heat extremes that IFS missed. Figure 12a is subsampled to assess the performance of HENS during instances of IFS busts. Figure 12b characterizes the relative occurrences of IFS busts and HENS busts using a confusion matrix. For the majority of cases (96%), both HENS and IFS capture the event: most of the time, they both have an ensemble member that is at least as large as the verification. However, in 3.5% of cases in summer 2023, IFS did not have a member with a sufficiently large 2m temperature value. For the vast majority of those cases, HENS did capture the true event. We note that there is a small portion of cases (0.32%) where HENS missed an event that IFS successfully captured. This is the topic of further research and could be due to cases where the HENS ensemble mean is biased.
As shown in Figure 12, many of the IFS busts occur around events in ERA5. For these events and other even rarer events, IFS cannot successfully sample the extremes, but HENS can.
@@ 5 Discussion and Conclusion
In total, HENS includes simulated days, or simulated years (7424 ensemble members 15 day rollout per ensemble member 92 forecast initialization days). With such a large sample size, we demonstrate the value-add of HENS: sampling events that are 4 standard deviations away from the ensemble mean, accurately estimating the tails of the ensemble, improving the skill of the best member, and reducing the likelihood of outliers. Based on its superior owCRPS scores and its narrow confidence intervals for extreme weather forecasts, HENS samples the tails of the forecast and observed distributions.
@@
For operational analysis of HENS, it is crucial to validate that the simulated weather events are trustworthy. We evaluate our HENS simulations on the medium-range timescale, up to the predictability limit of approximately 14 days. On these time scales, it is possible to directly compare the forecast with observations. The suite of NWP metrics provides insights into the realism of the ensemble forecast. If the ensemble members are exchangeable with each other and with the observations, the spread-error ratio (presented in Part I) should be close to 1. We show that SFNO-BVMC has a spread-error ratio that is close to 1, especially by a lead time of 10 days. Similarly, HENS has competitive scores on CRPS and twCRPS, which evaluate probabilistic forecasts against the true observed trajectories from ERA5. Additionally, Figure 12a shows that increasing the ensemble size allowed HENS to capture real events that were missed in the IFS ensemble. With its larger ensemble size, HENS reduced the number of instances where the verification dataset was greater than the ensemble max (Figure 12b). Increasing the ensemble size elevated the skill of the best ensemble member (Figure 7). With larger ensembles, at least one member better approximates the verification dataset. By increasing the sample size, the ensemble better simulates real extreme events, as measured through owCRPS. HENS improved its performance on these metrics focused on extremes (such as the outlier statistic, skill of the best member, and owCRPS), but it maintains its CRPS score. This increases trust in the ensemble predictions. To further validate whether the simulations realistically represent the dynamics of the Earth’s atmosphere, future research can validate the HENS predictions on idealized test cases (Hakim and Masanam, 2024; Mahesh et al., 2023; Mamalakis et al., 2022).
@@
An important future direction is to consider the effect of ensemble size in tandem with horizontal resolution. All our results here are based on 0.25-degree horizontal resolution, which is the resolution provided by ERA5. A key tradeoff in climate science is whether the compute budget is better spent on larger ensembles or finer resolution (Schneider et al., 2024). Because of their minimal computational cost, SFNO enables analysis of both questions. Future work is necessary to train ML ensembles to emulate kilometer-scale climate datasets (e.g. SCREAM (Caldwell et al., 2021)). This emulation can guide decisions on the tradeoff between ensemble size and horizontal resolution.
@@
For analyzing large simulations, an important technical frontier is the inline calculation of diagnostics. Traditionally, the simulation output is saved to disk during the forecast rollout, and then the saved data is loaded in a separate offline diagnostic analysis pipeline. On the other hand, inline diagnostics could be calculated during the ML ensemble generation itself. Offline diagnostic calculation is feasible for smaller ensembles that produce less data, but for HENS, it poses significant challenges due to I/O and disk storage space limits. Inline diagnostics could help alleviate the I/O and memory challenges associated with analyzing petabytes of simulation output. In particular, for climate analysis applications at scale, I/O can serve as the primary constraint for model analysis. In this manuscript, we save the entire simulation output and calculate our diagnostics offline, after the simulation is complete. We discuss the challenges and post-processing necessary for this in Section B. Inline diagnostics are an important frontier for future development with large data volumes from ML-based weather forecasts.
@@
Due to their cheap computational cost, ML-based forecasts present an opportunity to study the dynamics and statistics of extreme events. This opens the door for robust characterization of the drivers of extreme events, such as the heatwave drivers listed in Domeisen et al. (2022). With large dataset sizes, it becomes possible to apply causal analysis frameworks (Runge et al., 2019) to identify these drivers. For studying the statistics and drivers of rare events, a larger sample size is crucial. As a database of simulated weather events that is orders of magnitude larger than the observational record, HENS samples the tails of the weather forecast distributions and is a new resource for analyzing extreme weather.
@@
The code, datasets, and models are all stored at https://doi.org/10.5061/dryad.2rbnzs80n. The code is integrated with Zenodo: https://doi.org/10.5281/zenodo.14710345. We include the code to train SFNO, conduct ensemble inference with bred vectors and multiple checkpoints, and scoring and analysis code. We also open-source the model weights of the trained SFNO. See the README of the former DOI for information on how to use the codebase and for the permissive license associated with the code and data.
@@ Appendix B Post-processing Data to Improve Technical Analysis of the Huge Ensemble
Analysis of petabyte-scale data volumes is challenging and nontrivial. A key challenge relates to the storage of the ensemble in memory. For certain analysis below, we perform reductions on the ensemble dimension: at each initial time, lead time, latitude, and longitude, we calculate the ensemble mean, maximum, minimum, standard deviation, 99th percentile, bootstrap random sample, or CRPS. However, for a given initial time, HENS generates one file per ensemble member. This creates the following dimension ordering of the data in its storage in memory: (ensemble, lead time, latitude, longitude). For a given initial time, lead time, latitude, and longitude, the ensemble members are stored in separate locations in memory, not in a contiguous chunk. With our analysis workflow, this is suboptimal for performing ensemble reductions. It leads to prohibitive I/O read times to load each ensemble member and rearrange the data in memory such that the ensemble dimension can be reduced. Additionally, loading in the ensemble members in parallel and communicating via the Message Passing Interface (MPI) requires a significant number of concurrent processes and has prohibitive memory demands.
@@
Appendix C Effective of Ensemble Size on Reliability Diagrams and Spread Error Ratio
@@ Appendix E Example of IFS Busts and HENS Busts
Appendix F Statistical properties of the ensemble
F.1 Setup and notation
@@
F.2 Exchangeability
We first assess the extent to which there is a “signature” of the model checkpoints in the day-10 forecasts; i.e., are the output of the different model checkpoints interchangeable? This is commonly (in statistics) referred to as exchangeability. One way to assess the relative differences between model checkpoints is to propose a statistical model that compares the between-model (or inter-model) variability and the within-model (or intra-model) variability. If we re-index the HENS output to be , where indexes the individual model checkpoints and indexes the ensemble members for each checkpoint (yielding ), the statistical model assumes
@@ F.2 Exchangeability
F.3 Expected information gain
@@
F.3.1 Special case: Gaussian
@@
which can be solved with numerical integration. expected gain for Gaussian data is shown in Figure LABEL:Appdx_Fig_gausgauss_gain for values of ranging from to . As a sanity check, Figure 18 also shows a Monte Carlo estimate of the expected Gain for Gaussian data.
F.3.2 Information gain for HENS output
@@
F.4 Large-sample behavior of ensemble statistics
Next, we derive a theory for the large-sample behavior of seven ensemble statistics:
@@ F.4 Large-sample behavior of ensemble statistics
F.4.1 Analytic uncertainty in ensemble statistics
@@
F.4.2 Extreme value theory for “extreme” percentiles
@@
Central limit theorem for estimating percentiles with the generalized Pareto distribution. The Central limit theorem says that maximum likelihood estimates of the three parameters in Equation 12 have an asymptotically Normal distribution wherein the uncertainties scale with . Statistical theory allows us to derive the approximate standard error of an estimate of the percentile based on data points, denoted . Noting that the estimate is a function of as given in Equation 12, i.e.,
@@
F.4.3 Estimating ensemble statistics and uncertainties
To estimate the statistics of the ensemble and assess uncertainties as a function of ensemble size, we use Monte Carlo techniques similar to Section F.3.2. We estimate all statistics empirically and, for the 0.001 and 0.999 quantiles, additionally using extreme value theory. For an arbitrary statistic , the empirical calculation proceeds as follows: for and a given ,
@@ F.4.3 Estimating ensemble statistics and uncertainties
F.5 Aggregation
Lastly, we want to aggregate over all 92 days in the HENS simulation for each of the variables. However, there is clear seasonality in the ensemble mean and standard deviation for most (if not all) of the variables. To obviate these systematic differences, we summarize all quantities in terms of standard deviations. The expected information gain quantity is already in units of standard deviations, so we can simply average the expected information gain calculated for each initialization date. Furthermore, the exchangeability ratio in Section F.2 is also unitless and can hence be averaged over all initialization dates.

### v3

@@ 4.2 Confidence Intervals of Extreme Forecasts
In summer 2023, across all 10-day nonzero extreme forecasts at all grid cells, the confidence intervals are an order of magnitude narrower than those from small ensembles (Figure 10). This result is robust across different lead times: Figure 16 calculates the same confidence interval width at lead times of 4 days and 7 days. With less sampling uncertainty for extreme forecasts, there could be more informed disaster readiness and more targeted plans. However, even with narrower confidence intervals, we note that it is still of course possible to have uncertain forecasts. For instance, if half the ensemble members predict extreme weather, the chance of extreme would still be 50%. This probability is rooted in initial condition and model uncertainty. HENS does not alleviate these uncertainties; it reduces the sampling uncertainty of the forecast distribution. Therefore, with HENS, the 95 percent confidence interval around the 50% estimate would be narrower.
@@ 4.3 Missed Events in HENS and IFS
The outlier statistic measures the ability of an ensemble to capture the true verification value (Buizza and Palmer, 1998). They define this statistic as the proportion of cases in which the verification lies outside the bounds of the ensemble forecast. Using HENS, we calculate this statistic as a function of ensemble size. For each ensemble size, we take 100 bootstrap random samples from HENS. At each grid cell, if the ERA5 value lies within ninety-five percent of the bootstrapped ensembles, then the ensemble size is deemed satisfactory for capturing the event. Otherwise, the ERA5 value is classified as an outlier.
Figure 11 shows larger ensembles reduce the probability of an outlier. At a lead time of 10 days, the 7,424 member ensemble is satisfactory for representing almost 99% of the ERA5 values. The reduction in the outlier statistic is robust across lead times of 4 days, 7 days, and 10 days. The 10-day forecasts have a systematically lower outlier statistic than 7-day and 4-day forecasts because they have the best spread-error ratio (see Part 1). By taking 100 bootstrap random samples, we modify the calculation of the original statistic presented in Buizza and Palmer (1998). Their original statistic does not use bootstrap samples of the ensemble, so it does not account for sampling uncertainty. Our modification to account for sampling uncertainty tends to increase the outlier statistic. This modification requires 95% of all possible bootstrapped ensembles to capture the event, rather than just one ensemble. For a more direct comparison with the Buizza and Palmer (1998) method, we next calculate the proportion of cases in which the verification dataset is greater than the ensemble max (Figure 12). This is analogous to the warm side of their outlier statistic. In Figure 12, we do not consider ensemble sampling uncertainty, as in Figure 11. We only analyze warm outliers because summer 2023 was the hottest summer on record, and we focus on the ability of HENS to represent extreme heatwaves.
@@
Across all of summer 2023, Figure 12a shows that the bulk of the distribution lies above the 1-to-1 line, indicating that the HENS maximum successfully captures the heat extremes that IFS missed. Figure 12a only assesses the performance of HENS during instances of IFS busts. Figure 12b characterizes the relative occurrences of IFS busts and HENS busts using a confusion matrix. For the majority of cases (96%), both HENS and IFS capture the event: most of the time, they both have an ensemble member that is at least as large as the verification. However, in 3.5% of cases in summer 2023, IFS did not have a member with a sufficiently large 2m temperature value. For the vast majority of these IFS busts, HENS did capture the true event. We note that there is a small portion of cases (0.32%) where HENS missed an event that IFS successfully captured. This is the topic of further research and could be due to cases where the HENS ensemble mean is biased.
As shown in Figure 12, many of the IFS busts occur around events that have a Z score of approximately 2. For these events and other even rarer events, IFS cannot successfully sample the extremes. HENS can capture these events, yet it still maintains its CRPS (Figure 8), reliability (Figure 13), ensemble mean RMSE (Figure 14), ensemble spread (Figure 14), and spread-error ratio (Figure 15) of the corresponding 58-member SFNO-BVMC ensemble.
@@ 5 Discussion and Conclusion
In total, HENS includes simulated days, or simulated years (7424 ensemble members 15 day rollout per ensemble member 92 forecast initialization days). With such a large sample size, we demonstrate the value-add of HENS: sampling events that are 4 standard deviations away from the ensemble forecast mean, accurately estimating the tails of the ensemble, elevating the skill of the best member, and reducing the likelihood of outliers. Based on its superior owCRPS scores and its narrow confidence intervals for extreme weather forecasts, HENS samples the tails of the forecast and observed distributions.
@@
We evaluate our HENS simulations on the medium-range timescale, up to the predictability limit of approximately 14 days. On these time scales, it is possible to directly compare the forecast with observations. The suite of NWP metrics provides insights into the realism of the ensemble forecast. If the ensemble members are exchangeable with each other and with the observations, the spread-error ratio (presented in Part I) should be close to 1. We show that SFNO-BVMC has a spread-error ratio that is close to 1, especially by a lead time of 10 days. Similarly, HENS has competitive scores on CRPS and twCRPS, which evaluate probabilistic forecasts against the true observed trajectories from ERA5. Additionally, Figure 12a shows that increasing the ensemble size allowed HENS to capture real events that were missed in the IFS ensemble. With its larger ensemble size, HENS reduced the number of instances where the verification dataset was greater than the ensemble max (Figure 12b). Increasing the ensemble size increased the skill of the best ensemble member (Figure 7). By increasing the sample size, the ensemble better simulates real extreme events, as measured through owCRPS. To further validate whether the simulations realistically represent the dynamics of the Earth’s atmosphere, future research can validate the HENS predictions on idealized test cases (Hakim and Masanam, 2024; Mahesh et al., 2023; Mamalakis et al., 2022).
A fundamental constraint with simulating LLHIs is that these events are rare by definition, and there are limited observational samples with which to benchmark ensemble forecasting systems. With machine learning, this challenge is further complicated, since forty years of observations are reserved for training. Here, we demonstrate that huge ensembles of ML-based forecasting systems offer promising results for summer 2023. Future research is necessary to validate these ensemble systems on more LLHIs. In particular, the climate community can invest computational resources in creating large ensembles of physics-based simulations with high horizontal, vertical, and temporal resolution. These simulations would extend ML and LLHI science in multiple ways. In perfect model experiments, they can be used as additional out-of-sample simulations with which to validate ML weather prediction models. Alternatively, these simulations can be used to train the ML emulators, and all years of the observational record can be used as an out-of-sample validation set.
@@
An important future direction is to consider the effect of ensemble size in tandem with horizontal resolution. All our results here are based on 0.25-degree horizontal resolution, which is the resolution provided by ERA5. A key tradeoff in climate science is whether the compute budget is better spent on larger ensembles or finer resolution (Schneider et al., 2024). Because of their minimal computational cost, SFNO enables analysis of both options. Future work is necessary to train ML ensembles to emulate kilometer-scale climate datasets (e.g. SCREAM (Caldwell et al., 2021)). This emulation can guide decisions on the tradeoff between ensemble size and horizontal resolution.
@@
To analyze simulations with large data volumes, an important technical frontier is the inline calculation of diagnostics. Traditionally, the simulation output is saved to disk during the forecast rollout, and then the saved data is loaded into a separate offline diagnostic analysis pipeline. On the other hand, inline diagnostics could be calculated during the ML ensemble generation itself. Offline diagnostic calculation is feasible for smaller ensembles that produce less data, but for HENS, it poses significant challenges due to I/O and disk storage space limits. Inline diagnostics could help alleviate the I/O and memory challenges associated with analyzing petabytes of simulation output. In particular, for climate analysis applications at scale, I/O can serve as the primary constraint for model analysis. In this manuscript, we save the entire simulation output and calculate our diagnostics offline, after the simulation is complete. We discuss the challenges and post-processing necessary for this in Section B.
@@
Due to their cheap computational cost, ML-based forecasts present an opportunity to study the dynamics and statistics of extreme events. This opens the door for robust characterization of the drivers of extreme events, such as the heatwave drivers listed in Domeisen et al. (2022). With large dataset sizes, it becomes more feasible to apply causal analysis frameworks (Runge et al., 2019) to identify these drivers. For studying the statistics and drivers of rare events, a larger sample size is crucial. As a database of simulated weather events, HENS samples the tails of weather forecast distributions and is a new resource for analyzing extreme weather.
@@
The code, datasets, and models are all stored at https://doi.org/10.5061/dryad.2rbnzs80n, via DataDryad. The code is integrated with Zenodo at the prior DOI. We include the code to train SFNO, conduct ensemble inference with bred vectors and multiple checkpoints, and scoring and analysis code. We also open-source the model weights of the trained SFNO. See the README of the DOI for information on how to use the codebase and for the permissive license associated with the code and data. The code is available via the Lawrence Berkeley Lab BSD variant license, and the data is available with a CC0 license. For convenience, the webpage of our project is https://github.com/ankurmahesh/earth2mip-fork.
@@ Appendix B Post-processing Data to Improve Technical Analysis of the Huge Ensemble
Analysis of petabyte-scale data volumes is challenging and nontrivial. A key challenge relates to the storage of the ensemble in memory. For certain analysis below, we perform reductions on the ensemble dimension: for each initial time, lead time, latitude, and longitude, we calculate the ensemble mean, maximum, minimum, standard deviation, 99th percentile, bootstrap random sample, or CRPS. However, for a given initial time, HENS generates one file per ensemble member. This creates the following dimension ordering of the data in its storage in memory: (ensemble, lead time, latitude, longitude). For a given initial time, lead time, latitude, and longitude, the ensemble members are stored in separate locations in memory, not in a contiguous chunk. With our analysis workflow, this is suboptimal for performing ensemble reductions. It leads to prohibitive I/O read times to load each ensemble member and rearrange the data in memory such that the ensemble dimension can be reduced. Additionally, loading in the ensemble members in parallel and communicating via the Message Passing Interface (MPI) requires a significant number of concurrent processes and has prohibitive memory demands.
@@
Appendix C Effect of Ensemble Size on Reliability Diagrams and Spread Error Ratio
@@ Appendix E Example of IFS Busts and HENS Busts
Appendix F Assessing Robustness of HENS Across Lead Times and Variables
Appendix G Statistical properties of the ensemble
G.1 Setup and notation
@@
G.2 Exchangeability
We first assess the extent to which there is a “signature” of the model checkpoints in the day-10 forecasts, i.e., are the output of the different model checkpoints interchangeable? This is commonly (in statistics) referred to as exchangeability. One way to assess the relative differences between model checkpoints is to propose a statistical model that compares the between-model (or inter-model) variability and the within-model (or intra-model) variability. If we re-index the HENS output to be , where indexes the individual model checkpoints and indexes the ensemble members for each checkpoint (yielding ), the statistical model assumes
@@ G.2 Exchangeability
G.3 Expected information gain
@@
G.3.1 Special case: Gaussian
@@
which can be solved with numerical integration. A plot of the expected gain for Gaussian data is shown in Figure 23 for values of ranging from to . As a sanity check, Figure 23 also shows a Monte Carlo estimate of the expected Gain for Gaussian data.
G.3.2 Information gain for HENS output
@@
G.4 Large-sample behavior of ensemble statistics
We derive a theory for the large-sample behavior of seven ensemble statistics:
@@ G.4 Large-sample behavior of ensemble statistics
G.4.1 Analytic uncertainty in ensemble statistics
@@
G.4.2 Extreme value theory for “extreme” percentiles
@@
Central limit theorem for estimating percentiles with the generalized Pareto distribution. The Central limit theorem says that maximum likelihood estimates of the three parameters in Equation 12 have an asymptotically Normal distribution, wherein the uncertainties scale with . Statistical theory allows us to derive the approximate standard error of an estimate of the percentile based on data points, denoted . Noting that the estimate is a function of as given in Equation 12, i.e.,
@@
G.4.3 Estimating ensemble statistics and uncertainties
To estimate the statistics of the ensemble and assess uncertainties as a function of ensemble size, we use Monte Carlo techniques similar to Section G.3.2. We estimate all statistics empirically and, for the 0.001 and 0.999 quantiles, additionally using extreme value theory. For an arbitrary statistic , the empirical calculation proceeds as follows: for and a given ,
@@ G.4.3 Estimating ensemble statistics and uncertainties
G.5 Aggregation
Lastly, we want to aggregate over all 92 days in the HENS simulation for each of the variables. However, there is clear seasonality in the ensemble mean and standard deviation for most (if not all) of the variables. To obviate these systematic differences, we summarize all quantities in terms of standard deviations. The expected information gain quantity is already in units of standard deviations, so we can simply average the expected information gain calculated for each initialization date. Furthermore, the exchangeability ratio in Section G.2 is also unitless and can hence be averaged over all initialization dates.